OpenStudy (anonymous):

find the range of the function 1/1-sinx

OpenStudy (anonymous):

can't be 0

OpenStudy (mathmath333):

\(\large \text{use the fact that}\) \(\large\tt \begin{align} \color{black}{ -1\leq sinx \leq 1\\~\\~\\~\\ f(x)=\dfrac{1}{1-sinx}\\~\\ f(270^{\circ})=\dfrac{1}{1-(-1)}\\~\\ ~~~~~~~~~~~=\dfrac{1}{1-(-1)}\\~\\ ~~~~~~~~~~~=\dfrac{1}{2}\\~\\~\\~\\ f(x)=\dfrac{1}{1-sinx}\\~\\ f(90^{\circ})=\dfrac{1}{1-1}\\~\\ ~~~~~~~~~=\dfrac{1}{0}\\~\\ ~~~~~~~~~=\infty\\~\\~\\~\\ R_{f(x)}=[\dfrac{1}{2},\infty) }\end{align}\)